• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 12:05:39 on July 1,2010

    Entry number 329276

    keyword=timing adjustments / to-do

    We increased the widths of T1 and T3, that go to the coincidence
    circuit, to 30 nsec, per agreement with Bogdan.

    This way we are guaranteed to see coincidences with beam.
    After running with physics coincidences for awhile (e- pi+)
    we can judge how much to adjust the timing and then make the
    widths narrower: 10 nsec. It's a short access to do this.

    All the triggers were lined up at the trigger supervisor.

    Still to do (2 or 3 hours):

    1. Gate timing on R-HRS
    2. Setup new trigger = T1.and.T3.and.GasC-R

    ------------------

    It probably wasn't mentioned yet, but we replaced the faulty
    TDC in roc1 slot 11. This works now.